# DSL or Code? Evaluating the Quality of LLM-Generated Algebraic Specifications: A Case Study in Optimization at Kinaxis

### v1

@@ 1. Introduction
Model-driven engineering (MDE) has long been recognized for the conceptual abstraction and analytical rigour it brings to the development of software-intensive systems. Yet, outside sectors with certification or strict assurance mandates, the economics of building and maintaining models have often limited broader industrial uptake; many organizations perceive the costs as outweighing the benefits (Hutchinson et al., 2011; Whittle et al., 2013; Pourali and Atlee, 2018). Large language models (LLMs) are rapidly changing this cost calculus. By reducing the effort required to generate and modify structured artifacts from textual descriptions, LLMs enable domain experts to work primarily in natural language (NL) while still preserving the benefits of modelling, without incurring the high associated costs. This flexibility has led to growing interest in LLM-assisted derivation of models from text, e.g., (Ferrari et al., 2024; Yang et al., 2024; Chen et al., 2025; Khamsepour et al., 2025).
@@
The issue of representation for LLM-generated specifications is not unique to the industrial setting that motivates our work. Similar questions can arise in other domains, where one must decide whether to generate models in a specialized language, or instead produce code directly in a general-purpose programming language. Examples of such domains, among many others, include transformation (Jouault et al., 2008), system configuration (Dolstra et al., 2010; Labs, 2025), orchestration (Monsanto et al., 2012), testing (Cucumber, 2024), and rule checking (Maier et al., 2018). Under delivery pressure, teams may prefer “touching up” LLM-generated code (e.g., in Python) rather than revising a higher-level LLM-generated model that requires more substantial fixes. In other words, the potentially stronger out-of-the-box performance of LLMs in generating code from text makes code a tempting shortcut, creating the risk of a shift away from models to code even though models are now less costly to build through LLM assistance.
The central focus of this paper is whether LLM-generated DSL models can achieve parity with LLM-generated code in the qualities that matter most to practitioners: fidelity to stated intent and a low defect rate. To explore this question, and reflecting the industrial context of our collaborating partner, we focus on AMPL (A Mathematical Programming Language) — a high-level, algebraic modelling language designed for capturing and solving large-scale mathematical optimization problems, such as linear programming, mixed-integer programming, and nonlinear programming (AMPL, 2025; Fourer et al., 1990). AMPL has strong traction at our industry partner, thus providing a concrete setting for studying the trade-offs between DSL model generation and code generation using LLMs.
@@
(2) We empirically evaluate EXEOS, comparing AMPL models against Python code generated from identical inputs. Our evaluation is based on two datasets: a public corpus of optimization problems (Bertsimas and Tsitsiklis, 1997; Williams, 2013; Nace, 2020) and a set of real-world cases from Kinaxis in supply-chain planning. For each problem instance, we generate both an AMPL model and Python code, considering all combinations of the structuring step (present or absent) and iterative refinement (enabled or disabled) in EXEOS. This factorial design amounts to an ablation study, allowing us to isolate the effects of the target language (AMPL vs. Python), the structuring step, and repair. To ensure robustness, we experiment with two LLM families – GPT and Gemini – spanning both reasoning and instruction-following LLMs, and repeat each run several times, yielding over eleven thousand specifications. We assess executability by whether the specifications compile without errors, and correctness by comparing solutions to the ground truth using both exact matches and relative error.
Findings. The main findings from our evaluation are as follows: First, introducing a structuring step prior to generating a formal specification consistently improves quality, reducing compilation errors and yielding solutions more closely aligned with the intended optimization objectives. Second, iterative refinement further increases executability rates by enabling the automatic correction of specifications that initially failed at compile time or runtime. Third, generating Python code does not provide a systematic advantage over generating AMPL models. Specifically, across both datasets and LLM families, LLM-generated Python code shows no statistically significant improvement in correctness over AMPL. When EXEOS structures the NL problem description prior to specification generation and iteratively refines the generated specification, Python provides no executability benefit either: AMPL models execute more successfully on the public dataset and perform on par with Python on the industry cases. Indeed, AMPL models produced through the structure-generate-refine process generally outperform Python, especially when reasoning LLMs are used. This suggests that DSL model generation remains competitive, and in many cases even advantageous, when performed by sufficiently capable LLMs. Finally, our approach outperforms a baseline for automated optimization code generation without a separate data handling step as common in the literature (AhmadiTeshnizi et al., 2024), achieving higher executability and correctness in AMPL models.
Replication Package. All code, evaluation scripts, and experimental data for our public dataset are available online (Ayoughi, 2025).
@@ 2. Motivation
Both the Python program in Figure 3 and the AMPL model in Figure 3 are executable specifications of the optimization problem in Figure 1. The Python program is solver-specific, targeting the API of the Gurobi solver (Gurobi Optimization, LLC, 2025), whereas the AMPL model is solver-independent and can run with any compatible solver. For domain experts, the AMPL model offers clear advantages: in addition to being solver-independent, it is more compact and generally more natural than Python code. The difference is most evident in how objectives are expressed. In the Python program, lines 36–47 encode the revenue and holding-cost objectives; this requires knowledge of the Gurobi API and some programming workarounds. In particular, because Gurobi enforces a single global optimization sense (either maximization or minimization), all objectives must conform to that requirement. Thus, when the solver is set to maximize, the holding-cost objective (which should naturally be minimized) must be reformulated with a negative weight. Although valid, this approach makes the Python code harder to interpret or manually adjust. By contrast, the AMPL model on lines 22–27 expresses the objectives in a form much closer to the original problem description.
@@ 3. The AMPL Language
Mathematical optimization problems (optimization problems for short hereafter) formally model decision-making tasks, enabling one to determine the optimal values of related variables in order to achieve specific objectives. A Mathematical Programming Language (AMPL) (AMPL, 2025; Fourer et al., 1990) is a high-level modelling language designed for formulating and solving optimization problems. AMPL enables users to express optimization problems in a declarative, algebraic notation that closely resembles standard mathematical formulations. AMPL separates problem specification from data, allowing the same formulation to be applied to different datasets by updating only the data. AMPL also supports a broad range of solvers, enabling problems to be solved by different ones without altering their formulation. These flexible features have made AMPL suitable for use in both industry and academic research (AMPL, 2025; Fourer et al., 1990).
@@ 4. Our Approach
The approach has four steps: Step 1 structures the NL problem description and augments it with metadata. Step 2 processes the input data into a solver-ready data file. Step 3 generates a formal optimization specification, or regenerates it based on feedback from Step 4 if the solver fails due to compilation or runtime errors. In Step 4, the data file from Step 2 and the specification from Step 3 are used to compute a solution. Steps 1 and 3 require an LLM. Although different LLMs could be used, we employ a single instance for efficiency, given the sequential nature of the steps. We refer to this instance as LLM throughout the paper. The full list of prompts for EXEOS is available online (Ayoughi et al., 2025a). We next detail the steps of EXEOS.
Step 1. Structure and Add Metadata. This step identifies the main components of the optimization problem from the input description, organizes them into a structured NL representation, and extracts the metadata needed to build a mathematical formulation. Generally speaking, an optimization problem has four components (Bertsimas and Tsitsiklis, 1997; Williams, 2013; Nace, 2020): (1) objectives, which specify the quantities to be minimized or maximized. (2) parameters, which are known problem-specific constants. (3) variables, which represent decision points whose optimal values are determined by solving the problem. (4) constraints, which are expressed as algebraic relations between variables and parameters, restricting permissible variable values. For example, in the problem description of Figure 1, the objectives – maximizing revenue from product sales and minimizing inventory costs – are highlighted in pink. The parameters, in green, include product and resource types, resource requirements, current inventories, purchase costs, budget, inventory costs, and selling prices. The decision variables, in red, are the production quantities and additional resource purchases. Finally, the constraints, in brown, limit resource use to available and purchased units and restrict purchases to the budget.
@@
Step 4. Solve the optimization problem. This step uses an optimization solver such as Gurobi (Gurobi Optimization, LLC, 2025) or CPLEX (IBM, 2025), the formal specification from Step 3, and the solver-ready data from Step 2 to compute a solution. As shown in Figure 4, if attempting to solve the problem results in compilation or runtime errors, EXEOS initiates a refinement loop by sending the current specification and solver feedback to Step 3 (as previously explained) to regenerate the specification. This loop continues until the optimization problem is solved without errors or a predefined iteration limit is reached.
@@ 5.2. Baseline
To our knowledge, no prior work has addressed generating specifications for AMPL or other optimization DSLs. Existing work on optimization-specification generation mainly uses LLMs to derive Python programs from NL descriptions (AhmadiTeshnizi et al., 2024). Similar to EXEOS, this line of work applies a refinement loop when the generated specification fails to execute and further includes a step for structuring NL problem descriptions. However, because existing approaches are designed to process NL with embedded data values – as is common in academic exemplars – they lack a dedicated data-handling step and do not make parameter and variable metadata explicit. Having a step akin to Step 2 in EXEOS is nonetheless important for producing solver-ready data in large-scale industrial optimization. Motivated by this observation, our baseline adopts a variant of EXEOS that omits an explicit Step 2: data is extracted from the NL descriptions and inlined into the generated Python code.
@@ 5.3. Datasets
We use two datasets: (1) a public dataset, Public, containing 60 NL descriptions of optimization problems across five domains: facility location, network flow, scheduling, portfolio management, and energy optimization. These problem descriptions are drawn from three established sources in the optimization literature (Bertsimas and Tsitsiklis, 1997; Williams, 2013; Nace, 2020). Each problem has a data file and a ground-truth solution. The solver code and formal formulations for these problem descriptions are not publicly available, reducing the likelihood that LLMs encountered directly analyzable versions of these problems during training. (2) a proprietary dataset, Industry, from Kinaxis, with NL descriptions of six real-world supply-chain planning problems. Each problem includes a data file and a subject-matter-expert solution. Due to confidentiality, the Industry dataset cannot be released, and we are confident it was not used for LLM training. Table 2 provides summary statistics for the two datasets. The problem descriptions in Public include, on average, 1,018.12 characters and 171.61 tokens, while those in Industry contain, on average, 1,729 characters and 354 tokens. Thus, on average, the descriptions in Industry contain 1.7 $\times$ more characters and 2.06 $\times$ more tokens than those in Public.
@@
The Public and Industry datasets consist exclusively of single-objective optimization problems. This characteristic reflects both theoretical and practical considerations: multi-objective problems do not yield a single solution but rather a set of trade-offs (the Pareto set (Luke, 2013)), which requires additional criteria to reduce the ground truth to one solution. Widely used solvers such as Gurobi (Gurobi Optimization, LLC, 2025) and CPLEX (IBM, 2025) reinforce this property, as they do not enumerate the full Pareto frontier but instead return a single Pareto-optimal solution based on user-specified schemes such as weighted-sum aggregation or lexicographic prioritization. For this reason, optimization problems are typically formulated as single-objective ones, and cases that originally involve multiple objectives are reformulated with explicit prioritization to ensure unique ground-truth solutions, making results reproducible and deterministic. That said, EXEOS is agnostic to this factor and supports both multi-objective and single-objective formulations; for example, it can generate specifications for the illustrative multi-objective problem in Figure 1.
@@ 5.5. Implementation
Our implementation supports all the EXEOS variants in Table 1. It is written in Python 3.10 and built on top of LangChain (v0.2.8), which facilitates structured interactions with LLMs through its expression language (LCEL). For the AMPL variants, we use prompt templates that generate AMPL model (.mod) and data (.dat) files. For the Python variants, the templates produce code targeting the Gurobi optimization API, with input data captured in JSON format. Our complete implementation is available online (Ayoughi, 2025).
@@ 5.6. Experimental Procedure
We applied the eight EXEOS variants from Table 1 to the Public and Industry datasets using four LLMs: GPT-4o (OpenAI, 2024), (GPT) o4-mini (OpenAI, 2025), Gemini 1.5-Flash (Team et al., 2024), and Gemini 2.5-Pro (Comanici et al., 2025).All four LLMs were run with their default inference-time hyperparameters. Our GPT experiments were conducted through OpenAI’s API, and our Gemini experiments through Google Vertex AI, both under Kinaxis’ business subscriptions. GPT-4o and Gemini 1.5-Flash are instruction-following LLMs, while o4-mini and Gemini 2.5-Pro are reasoning LLMs. Considering both types enables us to address RQ4, which examines how instruction-following and reasoning LLMs differ in their ability to generate optimization specifications.
To mitigate random variation, each experiment was repeated five times. In total, we evaluated 66 optimization problems with eight EXEOS variants across four LLMs, repeating each experiment five times. This resulted in $66\times 8\times 4\times 5=\text{10,560}$ formal specification instances. The generation of these instances took 484 hours ($\approx$20 days). To solve all specification instances, whether in AMPL or Python, we used the Gurobi solver (Gurobi Optimization, LLC, 2025).
@@ 5.7. Results
To address our research questions using the results in Tables 3 and 4, we apply the following statistical tests: $Z$-test (Montgomery and Runger, 2019), Mann-Whitney $U$ test (Conover, 1999), and Vargha-Delaney effect size $\hat{A}_{12}$ (Vargha and Delaney, 2000). Specifically, we use the $Z$-test for our executability metric, execution success rate (Success). We report $Z$- and $p$-values at the 5% significance level, concluding that variant A (the first variant in the comparison) outperforms B (the second variant) if $Z>0$ and $p<0.05$. Otherwise, variant B outperforms A if $Z<0$ and $p<0.05$.
@@
To account for multiple comparisons, the $p$-values for both the $Z$ and $U$ tests are adjusted using the Benjamini-Hochberg procedure (Benjamini and Hochberg, 1995). Due to space constraints, statistical test results for RQ1–RQ4 are provided online (Ayoughi et al., 2025b), and we summarize them when answering these RQs. Results for RQ5 are reported in the paper.
@@
For average relative error, the result patterns diverge slightly: on the Public dataset, Ampl4 matches Python4 (0.48 vs. 0.49), while on the Industry dataset it shows a slightly higher error (0.29 vs. 0.21). Overall, statistical tests indicate that the average reduction in relative error of Python4 over Ampl4 is not significant (see (Ayoughi et al., 2025b)).

### v2

@@ 1. Introduction
Model-driven engineering (MDE) has long been recognized for the conceptual abstraction and analytical rigour it brings to the development of software-intensive systems. Yet, outside sectors with certification or strict assurance mandates, the economics of building and maintaining models have often limited broader industrial uptake; many organizations perceive the costs as outweighing the benefits (Hutchinson11; Whittle13; Pourali18). Large language models (LLMs) are rapidly changing this cost calculus. By reducing the effort required to generate and modify structured artifacts from textual descriptions, LLMs enable domain experts to work primarily in natural language (NL) while still preserving the benefits of modelling, without incurring the high associated costs. This flexibility has led to growing interest in LLM-assisted derivation of models from text, e.g., (FerrariAA24; YangCCMV24; Chen25; khamsepour2025impact).
@@
The issue of representation for LLM-generated specifications is not unique to the industrial setting that motivates our work. Similar questions can arise in other domains, where one must decide whether to generate models in a specialized language, or instead produce code directly in a general-purpose programming language. Examples of such domains, among many others, include transformation (ATL08), system configuration (Dolstra10; Puppet), orchestration (Monsanto12), testing (gherkin), and rule checking (datalog). Under delivery pressure, teams may prefer “touching up” LLM-generated code (e.g., in Python) rather than revising a higher-level LLM-generated model that requires more substantial fixes. In other words, the potentially stronger out-of-the-box performance of LLMs in generating code from text makes code a tempting shortcut, creating the risk of a shift away from models to code even though models are now less costly to build through LLM assistance.
The central focus of this paper is whether LLM-generated DSL models can achieve parity with LLM-generated code in the qualities that matter most to practitioners: fidelity to stated intent and a low defect rate. To explore this question, and reflecting the industrial context of our collaborating partner, we focus on AMPL (A Mathematical Programming Language) — a high-level, algebraic modelling language designed for capturing and solving large-scale mathematical optimization problems, such as linear programming, mixed-integer programming, and nonlinear programming (ampl; fourer1990modeling). AMPL has strong traction at our industry partner, thus providing a concrete setting for studying the trade-offs between DSL model generation and code generation using LLMs.
@@
(2) We empirically evaluate EXEOS, comparing AMPL models against Python code generated from identical inputs. Our evaluation is based on two datasets: a public corpus of optimization problems (Bertsimas24; williams2013model; Nace20) and a set of real-world cases from Kinaxis in supply-chain planning. For each problem instance, we generate both an AMPL model and Python code, considering all combinations of the structuring step (present or absent) and iterative refinement (enabled or disabled) in EXEOS. This factorial design amounts to an ablation study, allowing us to isolate the effects of the target language (AMPL vs. Python), the structuring step, and repair. To ensure robustness, we experiment with two LLM families – GPT and Gemini – spanning both reasoning and instruction-following LLMs, and repeat each run several times, yielding over eleven thousand specifications. We assess executability by whether the specifications compile without errors, and correctness by comparing solutions to the ground truth using both exact matches and relative error.
Findings. The main findings from our evaluation are as follows: First, introducing a structuring step prior to generating a formal specification consistently improves quality, reducing compilation errors and yielding solutions more closely aligned with the intended optimization objectives. Second, iterative refinement further increases executability rates by enabling the automatic correction of specifications that initially failed at compile time or runtime. Third, generating Python code does not provide a systematic advantage over generating AMPL models. Specifically, across both datasets and LLM families, LLM-generated Python code shows no statistically significant improvement in correctness over AMPL. When EXEOS structures the NL problem description prior to specification generation and iteratively refines the generated specification, Python provides no executability benefit either: AMPL models execute more successfully on the public dataset and perform on par with Python on the industry cases. Indeed, AMPL models produced through the structure-generate-refine process generally outperform Python, especially when reasoning LLMs are used. This suggests that DSL model generation remains competitive, and in many cases even advantageous, when performed by sufficiently capable LLMs. Finally, our approach outperforms a baseline for automated optimization code generation without a separate data handling step as common in the literature (AhmadiTeshniziG24), achieving higher executability and correctness in AMPL models.
Replication Package. All code, evaluation scripts, and experimental data for our public dataset are available online (neayoughi2025optigen).
@@ 2. Motivation
Both the Python program in Figure 3 and the AMPL model in Figure 3 are executable specifications of the optimization problem in Figure 1. The Python program is solver-specific, targeting the API of the Gurobi solver (gurobi), whereas the AMPL model is solver-independent and can run with any compatible solver. For domain experts, the AMPL model offers clear advantages: in addition to being solver-independent, it is more compact and generally more natural than Python code. The difference is most evident in how objectives are expressed. In the Python program, lines 36–47 encode the revenue and holding-cost objectives; this requires knowledge of the Gurobi API and some programming workarounds. In particular, because Gurobi enforces a single global optimization sense (either maximization or minimization), all objectives must conform to that requirement. Thus, when the solver is set to maximize, the holding-cost objective (which should naturally be minimized) must be reformulated with a negative weight. Although valid, this approach makes the Python code harder to interpret or manually adjust. By contrast, the AMPL model on lines 22–27 expresses the objectives in a form much closer to the original problem description.
@@ 3. The AMPL Language
Mathematical optimization problems (optimization problems for short hereafter) formally model decision-making tasks, enabling one to determine the optimal values of related variables in order to achieve specific objectives. A Mathematical Programming Language (AMPL) (ampl; fourer1990modeling) is a high-level modelling language designed for formulating and solving optimization problems. AMPL enables users to express optimization problems in a declarative, algebraic notation that closely resembles standard mathematical formulations. AMPL separates problem specification from data, allowing the same formulation to be applied to different datasets by updating only the data. AMPL also supports a broad range of solvers, enabling problems to be solved by different ones without altering their formulation. These flexible features have made AMPL suitable for use in both industry and academic research (ampl; fourer1990modeling).
@@ 4. Our Approach
The approach has four steps: Step 1 structures the NL problem description and augments it with metadata. Step 2 processes the input data into a solver-ready data file. Step 3 generates a formal optimization specification, or regenerates it based on feedback from Step 4 if the solver fails due to compilation or runtime errors. In Step 4, the data file from Step 2 and the specification from Step 3 are used to compute a solution. Steps 1 and 3 require an LLM. Although different LLMs could be used, we employ a single instance for efficiency, given the sequential nature of the steps. We refer to this instance as LLM throughout the paper. The full list of prompts for EXEOS is available online (optigen-supp-a). We next detail the steps of EXEOS.
Step 1. Structure and Add Metadata. This step identifies the main components of the optimization problem from the input description, organizes them into a structured NL representation, and extracts the metadata needed to build a mathematical formulation. Generally speaking, an optimization problem has four components (Bertsimas24; williams2013model; Nace20): (1) objectives, which specify the quantities to be minimized or maximized. (2) parameters, which are known problem-specific constants. (3) variables, which represent decision points whose optimal values are determined by solving the problem. (4) constraints, which are expressed as algebraic relations between variables and parameters, restricting permissible variable values. For example, in the problem description of Figure 1, the objectives – maximizing revenue from product sales and minimizing inventory costs – are highlighted in pink. The parameters, in green, include product and resource types, resource requirements, current inventories, purchase costs, budget, inventory costs, and selling prices. The decision variables, in red, are the production quantities and additional resource purchases. Finally, the constraints, in brown, limit resource use to available and purchased units and restrict purchases to the budget.
@@
Step 4. Solve the optimization problem. This step uses an optimization solver such as Gurobi (gurobi) or CPLEX (cplex), the formal specification from Step 3, and the solver-ready data from Step 2 to compute a solution. As shown in Figure 4, if attempting to solve the problem results in compilation or runtime errors, EXEOS initiates a refinement loop by sending the current specification and solver feedback to Step 3 (as previously explained) to regenerate the specification. This loop continues until the optimization problem is solved without errors or a predefined iteration limit is reached.
@@ 5.2. Baseline
To our knowledge, no prior work has addressed generating specifications for AMPL or other optimization DSLs. Existing work on optimization-specification generation mainly uses LLMs to derive Python programs from NL descriptions (AhmadiTeshniziG24). Similar to EXEOS, this line of work applies a refinement loop when the generated specification fails to execute and further includes a step for structuring NL problem descriptions. However, because existing approaches are designed to process NL with embedded data values – as is common in academic exemplars – they lack a dedicated data-handling step and do not make parameter and variable metadata explicit. Having a step akin to Step 2 in EXEOS is nonetheless important for producing solver-ready data in large-scale industrial optimization. Motivated by this observation, our baseline adopts a variant of EXEOS that omits an explicit Step 2: data is extracted from the NL descriptions and inlined into the generated Python code.
@@ 5.3. Datasets
We use two datasets: (1) a public dataset, Public, containing 60 NL descriptions of optimization problems across five domains: facility location, network flow, scheduling, portfolio management, and energy optimization. These problem descriptions are drawn from three established sources in the optimization literature (Bertsimas24; williams2013model; Nace20). Each problem has a data file and a ground-truth solution. The solver code and formal formulations for these problem descriptions are not publicly available, reducing the likelihood that LLMs encountered directly analyzable versions of these problems during training. (2) a proprietary dataset, Industry, from Kinaxis, with NL descriptions of six real-world supply-chain planning problems. Each problem includes a data file and a subject-matter-expert solution. Due to confidentiality, the Industry dataset cannot be released, and we are confident it was not used for LLM training. Table 2 provides summary statistics for the two datasets. The problem descriptions in Public include, on average, 1,018.12 characters and 171.61 tokens, while those in Industry contain, on average, 1,729 characters and 354 tokens. Thus, on average, the descriptions in Industry contain 1.7 $\times$ more characters and 2.06 $\times$ more tokens than those in Public.
@@
The Public and Industry datasets consist exclusively of single-objective optimization problems. This characteristic reflects both theoretical and practical considerations: multi-objective problems do not yield a single solution but rather a set of trade-offs (the Pareto set (metaheuristicsbook)), which requires additional criteria to reduce the ground truth to one solution. Widely used solvers such as Gurobi (gurobi) and CPLEX (cplex) reinforce this property, as they do not enumerate the full Pareto frontier but instead return a single Pareto-optimal solution based on user-specified schemes such as weighted-sum aggregation or lexicographic prioritization. For this reason, optimization problems are typically formulated as single-objective ones, and cases that originally involve multiple objectives are reformulated with explicit prioritization to ensure unique ground-truth solutions, making results reproducible and deterministic. That said, EXEOS is agnostic to this factor and supports both multi-objective and single-objective formulations; for example, it can generate specifications for the illustrative multi-objective problem in Figure 1.
@@ 5.5. Implementation
Our implementation supports all the EXEOS variants in Table 1. It is written in Python 3.10 and built on top of LangChain (v0.2.8), which facilitates structured interactions with LLMs through its expression language (LCEL). For the AMPL variants, we use prompt templates that generate AMPL model (.mod) and data (.dat) files. For the Python variants, the templates produce code targeting the Gurobi optimization API, with input data captured in JSON format. Our complete implementation is available online (neayoughi2025optigen).
@@ 5.6. Experimental Procedure
We applied the eight EXEOS variants from Table 1 to the Public and Industry datasets using four LLMs: GPT-4o (openai2024gpt4ocard), (GPT) o4-mini (openai2025introducing), Gemini 1.5-Flash (team2024gemini), and Gemini 2.5-Pro (comanici2025gemini).All four LLMs were run with their default inference-time hyperparameters. Our GPT experiments were conducted through OpenAI’s API, and our Gemini experiments through Google Vertex AI, both under Kinaxis’ business subscriptions. GPT-4o and Gemini 1.5-Flash are instruction-following LLMs, while o4-mini and Gemini 2.5-Pro are reasoning LLMs. Considering both types enables us to address RQ4, which examines how instruction-following and reasoning LLMs differ in their ability to generate optimization specifications.
To mitigate random variation, each experiment was repeated five times. In total, we evaluated 66 optimization problems with eight EXEOS variants across four LLMs, repeating each experiment five times. This resulted in $66\times 8\times 4\times 5=\text{10,560}$ formal specification instances. The generation of these instances took 484 hours ($\approx$20 days). To solve all specification instances, whether in AMPL or Python, we used the Gurobi solver (gurobi).
@@ 5.7. Results
To address our research questions using the results in Tables 3 and 4, we apply the following statistical tests: $Z$-test (montgomery2019applied), Mann-Whitney $U$ test (Mann-Whitney), and Vargha-Delaney effect size $\hat{A}_{12}$ (vargha2000critique). Specifically, we use the $Z$-test for our executability metric, execution success rate (Success). We report $Z$- and $p$-values at the 5% significance level, concluding that variant A (the first variant in the comparison) outperforms B (the second variant) if $Z>0$ and $p<0.05$. Otherwise, variant B outperforms A if $Z<0$ and $p<0.05$.
@@
Due to space constraints, statistical test results for RQ1–RQ4 are provided online (optigen-supp-b), and we summarize them when answering these RQs. Results for RQ5 are reported in the paper.
@@
For average relative error, the result patterns diverge slightly: on the Public dataset, Ampl4 matches Python4 (0.48 vs. 0.49), while on the Industry dataset it shows a slightly higher error (0.29 vs. 0.21). Overall, statistical tests indicate that the average reduction in relative error of Python4 over Ampl4 is not significant (see (optigen-supp-b)).